Find the 9th, 24th, and nth terms of the arithmetic sequence.
a, a + 6, a + 12, ...

Answer by mananth(16946)

a, a + 6, a + 12, ...
d= a+6-a=6
d=6
a9 = a1+8d
a9 = a1+48
a24 = a1+23d = a1+138
an = a1 (n-1)6 = a1+ 6n-6